OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x + 2y = 4 x – 5y = 11

OpenStudy (unklerhaukus):

x= 4-2y

OpenStudy (anonymous):

(2y + 4) – 5y = 11 x – 5(2y + 4) = 11 (–2y + 4) – 5y = 11 x – 5(–2y + 4) = 11

OpenStudy (anonymous):

those are the choices

OpenStudy (unklerhaukus):

looking at what i have done and substituting into the second equation using x = (-2y +4) and (x) - 5y =11 the anCwer should pop out at you

OpenStudy (anonymous):

? no it doesnt

OpenStudy (unklerhaukus):

the swap the terms that are bracketed